• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:07:57 on May 1, 2008

    Entry number 236044

    keyword=Beam Time Accounting
    Hall A

    04/30/08 1600-2400
    Experiment #: E04-007
    Shift Leader: B. Norum
    Shift Workers: Y. Qiang

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.500.380.260.100.140.680.932.99
    Beam Available
    but Not in Use
    0.000.200.420.700.820.520.280.042.98
    Beam Not Available
    or Unacceptable
    1.000.300.200.050.070.340.040.032.04
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.000.800.580.300.180.480.720.975.03
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.200.420.700.820.520.280.032.97
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.200.420.700.820.520.280.032.97
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00
    Bbite0.000.000.000.000.000.000.000.000.00

    Comments: Until approximately 21:30, almost every beam rf trip was accompanied by an HV failure in the BB, ostensibly due to beam spray as beam was restored. An extensive beam retune followed. Starting at 22:30, BB HV remained on through beam trips. BANU time almost completely due to time required to reset HV.

    Data Entry by: B. Norum